# Deep Bayes Factors

### v2

@@ 2.2 Deep Bayes Factors
The core idea behind Deep Bayes Factors is fairly simple and rests on the likelihood-ratio trick. To learn density ratios, this trick employs binary classification on two contrasting datasets, which are equipped with a binary label (Hastie et al., 2009; Sugiyama et al., 2012; Mohamed and Lakshminarayanan, 2016). The likelihood-ratio trick has been widely utilized in various contexts including generative adversarial networks (Goodfellow et al., 2020), noise-contrastive estimation (Gutmann and Hyvärinen, 2010), and mutual information estimation (Belghazi et al., 2018), among others. In the Bayesian literature, Kaji and Ročková (2022) used the likelihood-ratio trick for conditional likelihood ratio estimation inside the Metropolis-Hastings algorithm for posterior simulation. Here, we consider marginal likelihood ratio estimation for model comparisons.
@@
Recall that $P_{1}$ and $P_{2}$ correspond to the measures with marginal likelihoods $\pi(\cdot\,|\>M_{1})$ and $\pi(\cdot\,|\>M_{2})$, respectively. The optimal solution $D^{*}$ satisfies (Proposition 1 in (Goodfellow et al., 2020))
@@
Marginal likelihood ratio estimation is intimately linked to the task of estimating the $f$-divergence $d_{f}(P_{1}|P_{2})=\int\pi(\cdot\,|\>M_{1})f\left(\frac{\pi(\cdot\,|\>M_{2})}{\pi(\cdot\,|\>M_{2})}\right),$ where $f$ is a convex function satisfying $f(1)=0$. A prevalent approach for evaluating the divergence is through the Fenchel conjugate $f^{*}$, which yields a dual formula $d_{f}(P_{1}|P_{2})\geq\sup_{F\in\mathbb{T}}E_{1}[F]-E_{2}[f^{*}(F)]$ for a rich class functions $\mathbb{T}$. For instance, the original GAN Goodfellow et al. (2020) was extended to the f-GAN framework (Nowozin et al., 2016) by considering the $f$-distance between the original and fake distributions and optimizing over $\mathbb{T}$ using deep neural networks. It is known that choosing a suitable $f$ for specific tasks can lead to more stable optimization (Shannon, 2020). Because equality is achieved when $F^{*}=f^{\prime}\left(\frac{p_{1}}{p_{2}}\right)$, inverting $f^{\prime}$ in $F^{*}$ could yield the desired likelihood ratio. We adopted the classification formulation (3), acknowledging that other estimators could be obtained by choosing a suitable $f$ and inverting $f^{\prime}$. For a more comprehensive overview of likelihood ratio estimation through f-divergences, we refer the reader to Sugiyama et al. (2012).
In our demonstrations, we include all simulated samples $Y_{t}$ and $\tilde{Y}_{t}$ for $1\leq t\leq T$ as our training data for $\hat{D}_{T}$. Note that we do not require independence of the entries in $Y$, rendering this approach suitable for time series or otherwise correlated data. We could tailor the estimator to $Y_{0}^{(n)}$ by training the classifier on a screened-out ABC lookup table or using reinforcement-learning (sean).
@@
Throughout this paper, we primarily focus on deep learning classifiers $\mathcal{D}_{n}=\{d_{\phi}(\cdot):\phi\in\mathbb{R}^{d}\}$ parametrized by $\phi$. We describe the architecture choice in Section S3 and provide the code in Section S3.1. Utilizing stochastic optimization, we construct the training dataset on-the-fly. To train with as many samples as possible in a memory-efficient manner, we generate a new mini-batch of size $T/M$ from each model during every iteration, where $M$ represents the total number of iterations. A similar idea was used by shin2024generative, to avoid over-fitting since generating a new (mini)-batch for every single iteration maximizes the number of simulated data points used for the training. The described algorithm is in Algorithm 1. This algorithm produces a trained network $D^{(M)}$ after $M$ mini-batch iterations that we substitute for the estimated classifier $\hat{D}_{T}$. Algorithm 1 shows Bayes factor estimation evaluated on a single dataset $Y_{0}^{(n)}$. Note that after $\hat{D}_{T}$ has been trained, the Bayes factor estimator can be evaluated on any dataset $Y$.
@@
We apply the Deep Bayes Factor estimator (7) to three toy examples, for which closed-form Bayes factors exist. All of these examples assume an iid data setup. Data 1 explores classical Bayesian modeling with Negative Binomial and Poisson distributions (Bernardo and Smith, 2009). Define $M_{1}$ through $Y_{i}|p\sim NB(1,p)$, where $p\sim Beta(\alpha_{1},\beta_{1})$, and $M_{2}$ with $Y_{i}|\lambda\sim Pois(\lambda)$ where $\lambda\sim Gamma(\alpha_{2},\beta_{2})$. We set $\alpha_{1}=\beta_{1}=\alpha_{2}=\beta_{2}=1$. In this case, Robert et al. (2011) identified bias in ABC Bayes factors (2) when a summary statistic is used. Data 2 considers Gaussian mixture data with a Gaussian prior on the means. Namely, $M_{1}:Y_{i}\sim N(\mu_{11},2^{2})/2+N(\mu_{12},2^{2})/2$ with $\mu_{11}\sim N(2,1.5^{2}),~{}\mu_{12}\sim\frac{1}{2}N(-2,1.5^{2})$, and $M_{2}:Y_{i}\sim N(\mu_{2},2.5^{2})$ with $\mu_{2}\sim N(0,1)$. Data 3 presents a nested model test, distinguishing between a hierarchical Bayesian model $M_{1}:Y_{i}|\lambda\sim Exp(\lambda)$ with $\lambda\sim Gamma(2,2)$ and a fixed-parameter model $M_{2}:Y_{i}\sim Exp(3)$. We can conceptualize $M_{2}$ as a Bayesian hierarchical model with a Dirac prior $\delta(\lambda-3)$. To nicely visualize heatmaps of the results, we start with two observations, i.e. $n=2$, and present results for larger $n$ later in Section 6.1. The heatmaps zoom in onto values $Y=(Y_{1},Y_{2})^{\prime}$ that are most likely under either one of the two considered models (See Figure S16 in Section S5 for a broader range of $Y$). The ABC implementation is detailed in Section S2. In the effort to see the best ABC performance, we intentionally use the full data $Y$ as a summary statistic. Details on the implementations and deep learning architectures are in Section S3. The results are visualized in Figure 1, which shows high similarity between the true Bayes factors and the estimated ones on a large portion of the data domain that contains most of the support for both models. For the actual demonstration of the samples, see Figure S16 in Section S5.
@@ 2.3 Deep Bayes Factor Extensions
Deep Bayes Factors can be tailored to alternative Bayes factor formulations that leverage data splitting. Decomposing the data vector as $Y=(X,Z)^{\prime}$, where $X\in\mathcal{Y}^{n_{x}}$ and $Z\in\mathcal{Y}^{n_{z}}$, recall the conditional Bayes factor for $Z$ given $X$ (aka the partial Bayes factor PBF (Lempers, 1971; O’Hagan, 1995; Berger and Pericchi, 1996))
@@
where the last equality is due to $\pi(Z|X,M_{i})\pi(X|M_{i})=\pi(Z,X|M_{i})$. PBF’s were originally designed to handle non-informative (improper) priors or prior sensitivity by using the posterior $\pi_{k}(\theta_{k}|X)$ instead of the prior $\pi_{k}(\theta_{k})$ for $k=1,2$. The characterization (8) allows us to estimate PBF’s by deploying two Deep Bayes Factors (DeepBF’s). Another Bayes factor variant is the posterior Bayes factor, defined as $PBF_{1,2}(Z|X)$ in (8) where $X=Y$ and $Z=Y$. While one can estimate the posterior Bayes factor using deep learning via (8), it may not be preferred over alternatives because of its double use of the data. Another way of handling improper priors that is robust with respect to the choice of one particular split $(X,Z)$ is the intrinsic Bayes factor (IBF, Berger and Pericchi (1996)) which reflects all $K=\binom{n}{n_{x}}$ possible data subsets $Y(i)\in\mathcal{Y}^{n_{x}}$ of size $n_{x}$ and aggregates them through a geometric or arithmetic average. Seen as an actual Bayes factors for an intrinsic prior Berger and Pericchi (1996), the arithmetic IBF is defined as
@@
Figure 2: Examples of Bayes factor variants ($\log_{10}$ scale): PBF is a partial BF, PostBF is a posterior BF and GBF is a geometric variant of the intrinsic BF. The blanc area corresponds to numerical overflow (as elaborated on in Figure S17 and S16 (Section S5). The three datasets are from Example 1 with $n=2$.
@@ 3 Deep Bayes Factor Consistency
where $\mathcal{F}_{0}=\{\pi_{1}(\cdot\,|\>\theta_{1}):\theta_{1}\in\mathbb{R}^{d_{1}}\}$ and $\mathcal{F}_{1}=\{\pi_{2}(\cdot\,|\>\theta_{2}):\theta_{2}\in\mathbb{R}^{d_{2}}\}$. The likelihoods under each model are defined implicitly but are parametrized by finite-dimensional parameters. Below, we define the rate of Bayes factor divergence for the testing problem (10).
@@
Assumption 1 essentially states that the prior assigns enough mass around a shrinking neighborhood of $p^{*}$ under both considered models. Walker et al. (2004) characterized the Kullback-Leibler property as assigning a positive mass to a sufficiently large KL neighborhood and shows that Bayes factor always supports the model with a smaller KL gap, not necessarily requiring that the model arrived from either of the two models.
Before we establish model selection consistency of the Deep Bayes Factor estimator (7), we first focus on its consistency as a Bayes factor estimator. Theoretical results exist for classifiers obtained by minimizing the entropy loss (Kaji and Ročková, 2022; Kaji et al., 2020). The technique in Kaji and Ročková (2022) relies on properties of certain Hellinger-type discrepancy measures for classifiers. We can write the true Bayes factor as a functional of the “true” discriminator $D^{*}$ through the relationship $BF_{1,2}=D^{*}/(1-D^{*})$ in (4). In particular, under a variant of Assumption 1 in Kaji and Ročková (2022) (rephrased as Assumption 3 in the Supplement) the following statement holds for the classification-based estimator $\hat{D}_{T}$ defined in (5). For a non-negative sequence $\delta_{T,n}$ determining the quality of the training algorithm for the discriminator we have
@@
Note that the expectation in Theorem 1 is with respect to the marginal distribution $\pi(\cdot\,|\>M_{1})$ as opposed to the product measure $P^{\infty}_{*}$. Lemma 1 implies certain “consistency” of our estimator in the event that $\delta_{T,n}\rightarrow 0$ for any $n$ as $T\rightarrow\infty$, i.e. when the class of discriminators $\mathcal{D}_{n}$ is rich enough to be able to approximate $D^{*}$ sufficiently close. Our approximability/closeness assumption is stated implicitly inside Assumption 3 in the Supplement. For a class of classifiers $\mathcal{D}_{n}$, depending on the nature of $D^{*}$ (e.g. smoothness), we find the smallest $\delta_{T,n}$ for which this assumption is satisfied. Unlike with other approximability results for deep learning (see johannes and references therein), we do not have a precise characterization of $\delta_{T,n}$ in terms of the network width, depth and/or sparsity. We do acknowledge that there is still a gap between theory and practice of deep learning and that our method relies on a careful choice of the deep learning architecture which may not be entirely informed by our theory. We discuss practical aspects of picking the class of neural discriminators in Section S3 (Supplement). We use Lemma 1 as an intermediate step towards showing model selection consistency with $\log\widehat{BF}_{1,2}(Y_{0}^{(n)})$ shown in the following theorem.
@@
Theorem 1 shows that even if the Bayes factor estimator is inconsistent (i.e. $\delta_{T,n}$ does not vanish as $T\rightarrow\infty$), Deep Bayes Factors can still be consistent (in terms of model choice) as long as the true Bayes factors are consistent and $(n^{d_{1}/2}\vee n^{d_{2}/2})\delta_{T,n}$ is slower than the speed $1/\nu_{n}$ of Bayes factor divergence. Now, we review sufficient conditions under which the Bayes factor (and also our estimated Bayes factor) are, in fact, consistent for model selection when $p^{*}\in\mathcal{F}_{0}$. The case $p^{*}\in\mathcal{F}_{1}$ is symmetrical. These conditions essentially correspond to Tokdar et al. (2010) and Mcvinish et al. (2009).
@@
The actual Bayes factor ${BF}_{1,2}(Y_{0}^{(n)})$ converges to 00 in $P^{\infty}_{*}$-probability under Assumption 1(b) and Assumption 2 (according to Theorem 4.1 in Tokdar et al. (2010). A similar conclusion holds for $\widehat{BF}_{2,1}(Y_{0}^{(n)})$ within the context of Theorem 2 with the assumptions about priors $\pi_{1}$ and $\pi_{2}$ and models $M_{1}$ and $M_{2}$ reversed.
@@ 4 Generative AI for Bayesian Model Criticism
Because a model can be criticized without an explicit consideration of an alternative (Gelfand et al., 1992), it is difficult to entirely separate model selection from assessments of model adequacy. While our paper focuses on model selection, in this section we partially address how generative contrastive approaches could be used for model criticism. One of the key tools for Bayesian model criticism is the posterior predictive check (PPC) of Guttman (1967) whose premise is: “If my model is good, then its posterior predictive distribution will generate data that looks like my observations”. In PPC, the closeness is typically assessed through a diagnostic function, and a reference distribution has to be chosen to determine whether this diagnostic function indicates extreme values. An example diagnostic is the average squared distance to the posterior mean $\frac{1}{n}\sum_{i=1}^{n}(Y_{i}^{0}-E[Y_{i}\,|\>Y_{0}^{(n)}])^{2}$. A common criticism of such approaches is the double use of the data $Y_{0}$, for the reference distribution and for the evaluation of the diagnostic function. Avoiding the double of use of data, Bayarri and Berger (2000) propose the partial predictive check. Johnson (2007) proposed to use pivotal diagnostic to ensure the diagnostic and reference distributions are uncorrelated. Moran et al. (2019) considered population predictive checks based on data splitting. Below, we outline certain approaches which may serve as an alternative to these developments.
@@
During the training of $\hat{d}$, the observed data $Y_{0}^{(n)}$ is also used twice (first to simulate the fake data and then to train the classifier). However, this seems different from the usual double use of the data where $Y_{0}^{(n)}$ is used to construct a reference distribution and to evaluate the test statistic inside this distribution. Here, we can evaluate the test statistic $Z$ on the fake data $\tilde{Y}$, not the observed data $Y_{0}^{(n)}$. Instead of a “p-value” style measure of surprise, one can sample from the distribution $\pi(Z(\tilde{Y},\hat{d})\,|\>Y_{0}^{(n)})$ by evaluating $Z(\tilde{Y},\hat{d})$ on new plugged-in samples $\tilde{Y}\sim\pi(Y\,|\>Y_{0}^{(n)},M)$ from the posterior predictive distribution. Whether or not the model $M$ is good could be determined from the concentration of this distribution around $1/2$. Sampling from the parameter posterior using generative techniques (Wang and Ročková, 2022; Polson and Sokolov, 2023) makes sampling from $\pi(Z(\tilde{Y},\hat{d})\,|\>Y_{0}^{(n)},M)$ feasible. To alleviate systematic bias of a pre-trained classifier $\hat{d}$, one might need retrain $\hat{d}$ for each replication of $\tilde{Y}$. This strategy might be a new promising route towards implementing model criticism, i.e. compatibility of observed data with predictions from a model.
Figure 3: Model Criticism on Data 1. The empirical density (KDE) plot of $Z(\tilde{Y},\hat{d})$ values defined in (12). Red lines indicate the $2.5\%$ and $97.5\%$ quantiles of the simulated $Z(\tilde{Y},\hat{d})$ values. The blue line represents $0.5$.
Example 3 (Toy Example).
Using data setups from Example 1 and 2, we train Bayesian GAN posterior generators (Wang and Ročková, 2022). For each dataset, we generate one fixed $Y_{0}^{(n)}\sim M_{j}$ and then simulate $\tilde{Y}$ according to $\pi(\cdot\,|\>Y_{0}^{(n)},M_{k})$ for four model combinations $(j,k)\in\{1,2\}\times\{1,2\}$ and $n=128$. For $M_{2}$ of Data 3, we instead sample $\tilde{Y}\sim\pi(\cdot\,|\>M_{2})$ independently from $Y_{0}$. [^2] The model parameter is fixed as opposed to random with a prior. Additionally, we manually set ”outliers” as values $Y_{0}^{(n)}$ unlikely under the given model333 All outliers are in a form of $Y_{\rm out}=(\alpha,\alpha,\cdots,\alpha)\in\mathcal{Y}^{n}$ for a manually chosen $\alpha$. Dataset 1: $\alpha=18$, Dataset 2: $\alpha=10$, Dataset 3: $\alpha=2.5$.. In Figure 1, the first two dimensions of the outliers (red) and typical inliers (yellow) are marked as stars. To simulate $Z(\tilde{Y},\hat{d})$ values defined in (12), we train a simple logistic regression classifier, where the non-linear predictor is a quadratic function of a univariate input. The classifiers are optimized for every new draw of $\tilde{Y}$ and evaluated on a new draw of $\tilde{Y}$ to compute the draw $Z(\tilde{Y},\hat{d})$. We had 3 00030003\,0003 000 draws for each model.
The kernel density smoothed empirical distribution of $Z(\tilde{Y},\hat{d})$ is displayed in Figure 3 for Data 1, and in Figure S19 and S20 (Section S5) for Data 2 and 3, respectively. The consonant cases $\tilde{Y}\sim\pi(\cdot\,|\>Y_{0}^{(n)},M_{k})$ and $Y^{(n)}_{0}\sim\pi(\cdot\,|\>M_{k})$ for $k=1,2$ are illustrated on the diagonal of the Figure 3 (panels (a) and (e)). We can see that when the posterior predictive distribution matches the true generating distribution of $Y_{0}^{(n)}$, the $Z(\tilde{Y},\hat{d})$ values are concentrated near $0.5$. On the other hand, panels (b) and (d) show the dissonant cases where the true generating model and predictive model are mismatched. In these plots, the $Z(\tilde{Y},\hat{d})$ values are more spread out and the value $0.5$ is barely included within the the $2.5\%$ and $97.5\%$ sample quantiles. This indicates poor or borderline evidence for the considered model suggesting further model comparisons through Bayes factors. On the other hand, when $Y_{0}^{(n)}$ are implausible under the considered model, critiquing the suggested model becomes easier. This is evident from panels (c) and (f) where the kernel density plots of sampled values $Z(\tilde{Y},\hat{d})$ are far away from $0.5$. Such a scenario yields a more decisive rejection of the proposed model without the need for further Bayes factor model comparisons. It is important to note, however, Bayes factor values evaluated at observed data $Y^{(n)}_{0}$ implausible under both models may still strongly prefer one model over the other, despite the fact that predicted values from both modes are not compatible with $Y^{(n)}_{0}$. Therefore, it would seem that Bayes factor model comparisons could be enhanced with our Z-value approach which offers a careful preliminary examination of model adequacy before proceeding to model selection based on Bayes factors.

### v3

@@ 2.2 Deep Bayes Factors
The core idea behind Deep Bayes Factors is fairly simple and rests on the likelihood-ratio trick. To learn density ratios, this trick employs binary classification on two contrasting datasets, which are equipped with a binary label [31, 78, 54]. The likelihood-ratio trick has been widely utilized in various contexts including generative adversarial networks [26], noise-contrastive estimation [29], and mutual information estimation [6], among others. In the Bayesian literature, [41] used the likelihood-ratio trick for conditional likelihood ratio estimation inside the Metropolis-Hastings algorithm for posterior simulation. Here, we consider marginal likelihood ratio estimation for model comparisons.
@@
Recall that $P_{1}$ and $P_{2}$ correspond to the measures with marginal likelihoods $\pi(\cdot\,|\>M_{1})$ and $\pi(\cdot\,|\>M_{2})$, respectively. The optimal solution $D^{*}$ satisfies (Proposition 1 in [26])
@@
Marginal likelihood ratio estimation is intimately linked to the task of estimating the $f$-divergence $d_{f}(P_{1}|P_{2})=\int\pi(\cdot\,|\>M_{1})f\left(\frac{\pi(\cdot\,|\>M_{2})}{\pi(\cdot\,|\>M_{2})}\right),$ where $f$ is a convex function satisfying $f(1)=0$. A prevalent approach for evaluating the divergence is through the Fenchel conjugate $f^{*}$, which yields a dual formula $d_{f}(P_{1}|P_{2})\geq\sup_{F\in\mathbb{T}}E_{1}[F]-E_{2}[f^{*}(F)]$ for a rich class functions $\mathbb{T}$. For instance, the original GAN [26] was extended to the f-GAN framework [56] by considering the $f$-distance between the original and fake distributions and optimizing over $\mathbb{T}$ using deep neural networks. It is known that choosing a suitable $f$ for specific tasks can lead to more stable optimization [76]. Because equality is achieved when $F^{*}=f^{\prime}\left(\frac{p_{1}}{p_{2}}\right)$, inverting $f^{\prime}$ in $F^{*}$ could yield the desired likelihood ratio. We adopted the classification formulation (3), acknowledging that other estimators could be obtained by choosing a suitable $f$ and inverting $f^{\prime}$. For a more comprehensive overview of likelihood ratio estimation through f-divergences, we refer the reader to [78].
In our demonstrations, we include all simulated samples $Y_{t}$ and $\tilde{Y}_{t}$ for $1\leq t\leq T$ as our training data for $\hat{D}_{T}$. Note that we do not require independence of the entries in $Y$, rendering this approach suitable for time series or otherwise correlated data. We could tailor the estimator to $Y_{0}^{(n)}$ by training the classifier on a screened-out ABC lookup table or using reinforcement-learning [58].
@@
Throughout this paper, we primarily focus on deep learning classifiers $\mathcal{D}_{n}=\{d_{\phi}(\cdot):\phi\in\mathbb{R}^{d}\}$ parametrized by $\phi$. We describe the architecture choice in Section S3 and provide the code in Section S3.1. Utilizing stochastic optimization, we construct the training dataset on-the-fly. To train with as many samples as possible in a memory-efficient manner, we generate a new mini-batch of size $T/M$ from each model during every iteration, where $M$ represents the total number of iterations. A similar idea was used by [77], to avoid over-fitting since generating a new (mini)-batch for every single iteration maximizes the number of simulated data points used for the training. The described algorithm is in Algorithm 1. This algorithm produces a trained network $D^{(M)}$ after $M$ mini-batch iterations that we substitute for the estimated classifier $\hat{D}_{T}$. Algorithm 1 shows Bayes factor estimation evaluated on a single dataset $Y_{0}^{(n)}$. Note that after $\hat{D}_{T}$ has been trained, the Bayes factor estimator can be evaluated on any dataset $Y$.
@@
We apply the Deep Bayes Factor estimator (7) to three toy examples in an iid data setup, for which closed-form Bayes factors exist: Data 1) a comparison between Negative Binomial and Poisson distributions [11], Data 2) Gaussian mixture data with a Gaussian prior on the means, and Data 3) a nested model test distinguishing between hierarchical Bayesian models. More details are in Section S5.1. The results are visualized in Figure 1, which shows high similarity between the true Bayes factors and the estimated ones on a large portion of the data domain that contains most of the support for both models. For the actual demonstration of the samples, see Figure S18 in Section S7.
@@ 2.3 Deep Bayes Factor Extensions
Deep Bayes Factors can be tailored to alternative Bayes factor formulations that leverage data splitting. Decomposing the data vector as $Y=(X,Z)^{\prime}$, where $X\in\mathcal{Y}^{n_{x}}$ and $Z\in\mathcal{Y}^{n_{z}}$, recall the conditional Bayes factor for $Z$ given $X$ (aka the partial Bayes factor PBF [48, 57, 9])
@@
where the last equality is due to $\pi(Z|X,M_{i})\pi(X|M_{i})=\pi(Z,X|M_{i})$. PBF’s were originally designed to handle non-informative (improper) priors or prior sensitivity by using the posterior $\pi_{k}(\theta_{k}|X)$ instead of the prior $\pi_{k}(\theta_{k})$ for $k=1,2$. The characterization (8) allows us to estimate PBF’s by deploying two Deep Bayes Factors (DeepBF’s). Another Bayes factor variant is the posterior Bayes factor, defined as $PBF_{1,2}(Z|X)$ in (8) where $X=Y$ and $Z=Y$. While one can estimate the posterior Bayes factor using deep learning via (8), it may not be preferred over alternatives because of its double use of the data. Another way of handling improper priors that is robust with respect to the choice of one particular split $(X,Z)$ is the intrinsic Bayes factor (IBF, [9]) which reflects all $K=\binom{n}{n_{x}}$ possible data subsets $Y(i)\in\mathcal{Y}^{n_{x}}$ of size $n_{x}$ and aggregates them through a geometric or arithmetic average. Seen as an actual Bayes factors for an intrinsic prior [9], the arithmetic IBF is defined as
@@
Figure 2: Examples of Bayes factor variants ($\log_{10}$ scale): PBF is a partial BF, PostBF is a posterior BF and GBF is a geometric variant of the intrinsic BF. The blanc area corresponds to numerical overflow (as elaborated on in Figure S19 and S18 (Section S7). The three datasets are from Example 1 with $n=2$.
@@ 3 Deep Bayes Factor Consistency
where $\mathcal{F}_{0}=\{\pi_{1}(\cdot\,|\>\theta_{1}):\theta_{1}\in\Theta_{1}\}$ and $\mathcal{F}_{1}=\{\pi_{2}(\cdot\,|\>\theta_{2}):\theta_{2}\in\Theta_{2}\}$. The likelihoods under each model are defined implicitly but are parametrized by finite-dimensional parameters. Below, we define the rate of Bayes factor divergence for the testing problem (10).
@@
Assumption 1 essentially states that the prior assigns enough mass around a shrinking neighborhood of $p^{*}$ under both considered models. [88] characterized the Kullback-Leibler property as assigning a positive mass to a sufficiently large KL neighborhood and shows that Bayes factor always supports the model with a smaller KL gap, not necessarily requiring that the model arrived from either of the two models.
Before we establish model selection consistency of the Deep Bayes Factor estimator (7), we first focus on its consistency as a Bayes factor estimator. Theoretical results exist for classifiers obtained by minimizing the entropy loss [41, 42]. The technique in [41] relies on properties of certain Hellinger-type discrepancy measures for classifiers. We can write the true Bayes factor as a functional of the “true” discriminator $D^{*}$ through the relationship $BF_{1,2}=D^{*}/(1-D^{*})$ in (4). In particular, under a variant of Assumption 1 in [41] (rephrased as Assumption 3 in the Supplement) the following statement holds for the classification-based estimator $\hat{D}_{T}$ defined in (5). For a non-negative sequence $\delta_{T,n}$ determining the quality of the training algorithm for the discriminator we have
@@
Note that the expectation in Lemma 1 is with respect to the marginal distribution $\pi(\cdot\,|\>M_{1})$ as opposed to the product measure $P^{\infty}_{*}$. Lemma 1 implies certain “consistency” of our estimator in the event that $\delta_{T,n}\rightarrow 0$ for any $n$ as $T\rightarrow\infty$, i.e. when the class of discriminators $\mathcal{D}_{n}$ is rich enough to be able to approximate $D^{*}$ sufficiently close. Our approximability/closeness assumption is stated implicitly inside Assumption 3 in the Supplement. For a class of classifiers $\mathcal{D}_{n}$, depending on the nature of $D^{*}$ (e.g. smoothness), we find the smallest $\delta_{T,n}$ for which this assumption is satisfied. Unlike with other approximability results for deep learning (see [73] and references therein), we do not have a precise characterization of $\delta_{T,n}$ in terms of the network width, depth and/or sparsity. We do acknowledge that there is still a gap between theory and practice of deep learning and that our method relies on a careful choice of the deep learning architecture which may not be entirely informed by our theory. We discuss practical aspects of picking the class of neural discriminators in Section S3 (Supplement). We use Lemma 1 as an intermediate step towards showing model selection consistency with $\log\widehat{BF}_{1,2}(Y_{0}^{(n)})$ shown in the following theorem.
@@
Theorem 1 shows that even if the Bayes factor estimator is inconsistent (i.e. $\delta_{T,n}$ does not vanish as $T\rightarrow\infty$), Deep Bayes Factors can still be consistent (in terms of model choice) as long as the true Bayes factors are consistent and $(n^{d_{1}/2}\vee n^{d_{2}/2})\delta_{T,n}$ is slower than the speed $1/\nu_{n}$ of Bayes factor divergence. Now, we review sufficient conditions under which the Bayes factor (and also our estimated Bayes factor) are, in fact, consistent for model selection when $p^{*}\in\mathcal{F}_{0}$. The case $p^{*}\in\mathcal{F}_{1}$ is symmetrical. These conditions essentially correspond to [83] and [50].
@@
The actual Bayes factor ${BF}_{1,2}(Y_{0}^{(n)})$ converges to 00 in $P^{\infty}_{*}$-probability under Assumption 1(b) and Assumption 2 (according to Theorem 4.1 in [83]. A similar conclusion holds for $\widehat{BF}_{2,1}(Y_{0}^{(n)})$ within the context of Theorem 2 with the assumptions about priors $\pi_{1}$ and $\pi_{2}$ and models $M_{1}$ and $M_{2}$ reversed.
@@ 4 Generative AI for Bayesian Model Criticism
Because a model can be criticized without an explicit consideration of an alternative [21], it is difficult to entirely separate model selection from assessments of model adequacy. While our paper focuses on model selection, in this section we partially address how generative contrastive approaches could be used for model criticism. One of the key tools for Bayesian model criticism is the posterior predictive check (PPC) of [30] whose premise is: “If my model is good, then its posterior predictive distribution will generate data that looks like my observations”. In PPC, the closeness is typically assessed through a diagnostic function, and a reference distribution has to be chosen to determine whether this diagnostic function indicates extreme values. An example diagnostic is the average squared distance to the posterior mean $\frac{1}{n}\sum_{i=1}^{n}(Y_{i}^{0}-E[Y_{i}\,|\>Y_{0}^{(n)}])^{2}$. A common criticism of such approaches is the double use of the data $Y_{0}$, for the reference distribution and for the evaluation of the diagnostic function. Avoiding the double of use of data, [4] propose the partial predictive check. [40] proposed to use pivotal diagnostic to ensure the diagnostic and reference distributions are uncorrelated. [55] considered population predictive checks based on data splitting. Below, we outline certain approaches which may serve as an alternative to these developments.
@@
During the training of $\hat{d}$, the observed data $Y_{0}^{(n)}$ is also used twice (first to simulate the fake data and then to train the classifier). However, this seems different from the usual double use of the data where $Y_{0}^{(n)}$ is used to construct a reference distribution and to evaluate the test statistic inside this distribution. Here, we can evaluate the test statistic $Z$ on the fake data $\tilde{Y}$, not the observed data $Y_{0}^{(n)}$. Instead of a “p-value” style measure of surprise, one can sample from the distribution $\pi(Z(\tilde{Y},\hat{d})\,|\>Y_{0}^{(n)})$ by evaluating $Z(\tilde{Y},\hat{d})$ on new plugged-in samples $\tilde{Y}\sim\pi(Y\,|\>Y_{0}^{(n)},M)$ from the posterior predictive distribution. Whether or not the model $M$ is good could be determined from the concentration of this distribution around $1/2$. Sampling from the parameter posterior using generative techniques [89, 59] makes sampling from $\pi(Z(\tilde{Y},\hat{d})\,|\>Y_{0}^{(n)},M)$ feasible. To alleviate systematic bias of a pre-trained classifier $\hat{d}$, one might need retrain $\hat{d}$ for each replication of $\tilde{Y}$. This strategy might be a new promising route towards implementing model criticism, i.e. compatibility of observed data with predictions from a model. An example of this idea is presented in Section S5.2 in the supplement.